AO Com Work: 29 October 2010 UT

Observer: DMiller SEsposito, MXompero, LBusoni, APuglisi, FQuiros
Telescope Operator: SAllanson (LBTO)
Telescope Support: JLittle (LBTO)
Instrument Support: (LBTO)


A lot of AO data collected. No problems to speak of. Good weather and 0.8 to 1.3" seeing



-- DougMiller - 27 Oct 2010
Topic revision: r2 - 29 Oct 2010, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback